# Simultaneous Inference in Multiple Matrix-Variate Graphs for High-Dimensional Neural Recordings

### v1

@@ 3.2.1 Single Edge Test
By leveraging information across $m$ sessions using group lasso (Eq. 2), we not only improve the efficiency of spatial partial correlation estimates but also enhance the power of tests for detecting significant associations. We first consider testing a single edge: for a pair $(i,j)$, the null hypothesis is
@@
Given the group sparsity structure, we construct a test statistic by aggregating the partial correlation estimates $\widehat{\rho}^{(\mathcal{S},l)}_{ij}$ across all sessions, assuming the sign of the associations remains consistent. Thus, the sum of estimates approximates $\sum_{l=1}^{m}\lvert\rho^{(\mathcal{S},l)}_{ij}\rvert$, which equals zero under the null. To account for varying sample sizes across sessions, we weight the estimates by $\sqrt{n_{l}p}$ and define the test statistic as $\widehat{T}_{ij}:=\frac{1}{\sqrt{m}}\sum_{l=1}^{m}\sqrt{n_{l}p}\,\widehat{\rho}^{(\mathcal{S},l)}_{ij}.$ In Proposition 4.2, we show that $\widehat{T}_{ij}$ asymptotically follows a normal distribution with mean $T_{ij}:=\frac{1}{\sqrt{m}}\sum_{l=1}^{m}\sqrt{n_{l}p}\,\rho^{(\mathcal{S},l)}_{ij},$ so $\widehat{T}_{ij}$ should be significantly greater than zero under the alternative hypothesis. Once the asymptotic variance is consistently estimated, constructing a confidence interval for $T_{ij}$ and the corresponding $p$-value for $H_{0,ij}$ is straightforward. However, we shift our focus to the more challenging multiple edge test, treating the single edge test as a trivial special case.
@@
More generally, additional sign information on elements of the alternative $\rho^{(\mathcal{S},l)}_{ij}$ may be available. With this additional knowledge, we present a test statistic based on a linear combination of those $\widehat{\rho}^{(\mathcal{S},l)}_{ij}$ for $l=1,\dots,m$, which is closely related to its $\ell_{1}$ norm. More specifically, with an edge-specific sign vector $\sigma_{ij}:=(\sigma^{(1)}_{ij},\dots,\sigma^{(m)}_{ij})^{\top}\in\{-1,1\}^{m}$, we replace $\widehat{T}_{ij}$ with the following sign-addressed test statistic $\widehat{T}_{ij,\sigma}:=\frac{1}{\sqrt{m}}\sum_{l=1}^{m}\sigma^{(l)}_{ij}\sqrt{n_{l}p}\widehat{\rho}^{(\mathcal{S},l)}_{ij}$. The normal approximation we establish in Section 4 also applies to the sign-addressed test statistic.
@@ 3.2.2 Simultaneous Test
In the multiple edge test scenario, we aim to test whether there are no edges at all in a user-specified edge set $E$, which corresponds to the following null hypothesis:
@@
When $E$ consists of a single edge $(i,j)$, $H_{0,E}$ reduces to the single-edge null hypothesis $H_{0,ij}$ in Eq. 5. While multiple testing techniques, such as Bonferroni correction, can be applied to extend single-edge tests to multiple-edge tests, these methods are often overly conservative. Moreover, in neuroscience applications, the edge set $E$ typically represents connections between different brain areas, and its cardinality can grow up to $q^{2}$. As a result, even if we establish the asymptotic normality of a single $\widehat{T}_{ij}$, traditional multiple testing methods may not be valid in high-dimensional settings where $\lvert E\rvert$ grows rapidly.
@@
The key idea leverages high-dimensional central limit theory (e.g., Chernozhukov et al.,, 2013): although the full vector $\widehat{T}_{E}$ may not be asymptotically normal as $\lvert E\rvert$ increases, the supremum norm $\lVert\widehat{T}_{E}-T_{E}\rVert_{\infty}$ exhibits the same limiting behavior as $\lVert Z\rVert_{\infty}$, where $Z$ is a centered normal random vector with the same covariance as the asymptotic covariance of $\widehat{T}_{E}$. Specifically, this covariance is given by the matrix of asymptotic covariances between $\widehat{T}_{i_{1}j_{1}}$ and $\widehat{T}_{i_{2}j_{2}}$ for $(i_{1},j_{1}),(i_{2},j_{2})\in E$, expressed as:
$\begin{split}&S_{(i_{1},j_{1}),(i_{2},j_{2})}:=\\ &\sum_{l=1}^{m}\frac{\lVert\Sigma^{(\mathcal{T},l)}\rVert_{F}^{2}}{mp}{\scriptstyle\left[\begin{split}&\rho^{(\mathcal{S},l)}_{i_{1}i_{2}}\rho^{(\mathcal{S},l)}_{j_{1}j_{2}}+\rho^{(\mathcal{S},l)}_{i_{1}j_{2}}\rho^{(\mathcal{S},l)}_{i_{2}j_{1}}+\frac{1}{2}\rho^{(\mathcal{S},l)}_{i_{1}j_{1}}\rho^{(\mathcal{S},l)}_{i_{2}j_{2}}\Big{(}\rho^{(\mathcal{S},l)2}_{i_{1}i_{2}}+\rho^{(\mathcal{S},l)2}_{j_{1}j_{2}}+\rho^{(\mathcal{S},l)2}_{i_{1}j_{2}}+\rho^{(\mathcal{S},l)2}_{i_{2}j_{1}}\Big{)}\\ &-\rho^{(\mathcal{S},l)}_{i_{1}i_{2}}\rho^{(\mathcal{S},l)}_{i_{2}j_{2}}\rho^{(\mathcal{S},l)}_{i_{2}j_{1}}-\rho^{(\mathcal{S},l)}_{i_{1}i_{2}}\rho^{(\mathcal{S},l)}_{i_{1}j_{1}}\rho^{(\mathcal{S},l)}_{i_{1}j_{2}}-\rho^{(\mathcal{S},l)}_{j_{1}j_{2}}\rho^{(\mathcal{S},l)}_{i_{2}j_{2}}\rho^{(\mathcal{S},l)}_{i_{1}j_{2}}-\rho^{(\mathcal{S},l)}_{j_{1}j_{2}}\rho^{(\mathcal{S},l)}_{i_{2}j_{1}}\rho^{(\mathcal{S},l)}_{i_{1}j_{1}}\end{split}\right]}.\end{split}$ (7)
We approximate the distribution of $\lVert Z\rVert_{\infty}$ using a parametric bootstrap based on the plug-in estimator $\widehat{S}_{EE}$ of the asymptotic covariance. By generating bootstrap samples from $\widehat{Z}\sim\operatorname{N}(0,\widehat{S}_{EE})$, we construct a $(1-\alpha)$ confidence region:
$\mathcal{C}_{E}(1-\alpha):=\left\{T_{E}:\|\widehat{T}_{E}-T_{E}\|_{\infty}=\max_{(i,j)\in E}\left\lvert\frac{1}{\sqrt{m}}\sum_{l=1}^{m}\sqrt{n_{l}p}(\widehat{\rho}^{(\mathcal{S},l)}_{ij}-\rho^{(\mathcal{S},l)}_{ij})\right\rvert\leq\widehat{q}_{\lVert\widehat{Z}\rVert_{\infty},1-\alpha}\right\},$ (8)
where $\widehat{q}_{\lVert\widehat{Z}\rVert_{\infty},1-\alpha}$ is the bootstrap $(1-\alpha)$ quantile of $\lVert\widehat{Z}\rVert_{\infty}$. The null hypothesis $H_{0,E}$ is rejected if $\mathbf{0}\notin\mathcal{C}_{E}(1-\alpha)$. This confidence region can also be extended to $c$-level tests (Qiu and Zhou,, 2020). The coverage of this confidence region is studied in Theorem 4.3, with a power analysis provided in Theorem 4.4. We further discuss how borrowing information across sessions enhances testing power in Remark 4.5.
@@ 3.3 Estimation of Temporal Covariance Matrix
The plug-in estimator $\widehat{S}_{EE}$ of the asymptotic covariance (Eq. 7) requires the Frobenius norm of the temporal covariance matrix, $\Sigma^{(\mathcal{T},l)}$ for each session. We propose to estimate $\Sigma^{(\mathcal{T},l)}$ based on a modified Cholesky decomposition of its inverse, the temporal precision matrix $\Omega^{(\mathcal{T},l)}$ (Bickel and Levina,, 2008; Liu and Ren,, 2020). The Cholesky decomposition reveals the natural auto-regressive relationship between the signal at a particular time point and the past signals for spatiotemporal data. Suppose that the modified Cholesky decomposition of $\Omega^{(\mathcal{T},l)}$ is $\Omega^{(\mathcal{T},l)}=L^{(\mathcal{T},l)}D^{(\mathcal{T},l)}L^{(\mathcal{T},l)\top},$ where $D^{(\mathcal{T},l)}$ is a diagonal matrix, and $L^{(\mathcal{T},l)}$ is a lower triangular matrix with diagonal entries equal to $1$. Let $\beta^{(\mathcal{T},l)}:=I-L^{(\mathcal{T},l)\top}$ and $\Phi^{(\mathcal{T},l)}=\frac{\operatorname{tr}(\Sigma^{(\mathcal{S},l)})}{q}(D^{(\mathcal{T},l)})^{-1}$ so that the Cholesky decomposition of $\Omega^{(\mathcal{T},l)}$ can be rewritten by
$\Omega^{(\mathcal{T},l)}=\frac{\operatorname{tr}(\Sigma^{(\mathcal{S},l)})}{q}(I-\beta^{(\mathcal{T},l)})^{\top}(\Phi^{(\mathcal{T},l)})^{-1}(I-\beta^{(\mathcal{T},l)}).$ (9)
@@
Motivated by our working example, where temporal alignment of neural recordings is not guaranteed due to response latencies to stimuli (Ventura,, 2004), we do not assume that $\Sigma^{(\mathcal{T},l)}$ is the same or even similar across sessions. Instead, we impose a weaker bandable assumption on the modified Cholesky decomposition for each session. In physiological time-series signals, the dependence between time points naturally decays as the time lag increases, so a reasonable assumption is that $\beta^{(\mathcal{T},l)}_{st}$ approaches zero as $t-s\rightarrow\infty$. By applying bandable assumptions, we ignore weak dependencies between distant time points, achieving a bias-variance trade-off in the estimation of $\beta^{(\mathcal{T},l)}$.
@@
We estimate $\beta^{(\mathcal{T},l)}$ following the procedure in Liu and Ren, (2020). Specifically, we treat the observed data $X^{(k,l)}$ at trial $k$ and session $l$ as $q$ vector-variate “column samples” and fit the following linear regression model at each time point $t$:
@@
$\widehat{\beta}^{(\mathcal{T},l)}_{\cdot,t}:=\operatorname*{\arg\!\min}_{b\in\mathbb{R}^{p}}\frac{1}{2n_{l}q}\sum_{k=1}^{n_{l}}\lVert X^{(k,l)}_{t,\cdot}-X^{(k,l)\top}b\rVert_{2}^{2}$ (10)
@@
with resepct to $b_{s}=0$ where $s<t-h_{l}$ or $s\geq t$, where the bandwidth $h_{l}$ can be either user-specified or data-driven. Unlike the spatial case, the estimation of the temporal covariance matrix is performed individually for each session. Having the regression coefficients estimated, we estimate $\Phi^{(\mathcal{T},l)}$ by
@@
$\widehat{\Phi}^{(\mathcal{T},l)}_{tt}=\frac{1}{n_{l}q}\sum_{k=1}^{n_{l}}\left\lVert X^{(k,l)}_{t,\cdot}-X^{(k,l)\top}\beta^{(\mathcal{T},l)}_{\cdot,t}\right\rVert_{2}^{2}.$ (11)
@@
For technical issues, we truncate the eigenvalues of $I-\beta^{(\mathcal{T},l)}$, following the approach in Liu and Ren, (2020). For a square matrix $A$, let $P_{\eta}(A):=U~{}\max\{\min\{\Lambda,\eta\},\eta^{-1}\}~{}V^{\top}$, where $A$ has a singular value decomposition $A=U\Lambda V^{\top}$ and the $\min$ and $\max$ above are element-wise operations. Let $\overline{\Omega}^{(\mathcal{T},l)}$ be the precursor estimator of $\Omega^{(\mathcal{T},l)}$ given by
$\overline{\Omega}^{(\mathcal{T},l)}=P_{\eta}(I-\widehat{\beta}^{(\mathcal{T},l)})^{\top}(\widehat{\Phi}^{(\mathcal{T},l)})^{-1}P_{\eta}(I-\widehat{\beta}^{(\mathcal{T},l)}).$ (12)
@@
In the end, due to the identifiability constraint $\operatorname{tr}(\Sigma^{(\mathcal{T},l)})=p$, we propose our estimators of $\Sigma^{(\mathcal{T},l)}$ and $\Omega^{(\mathcal{T},l)}$ as
@@
$\widehat{\Sigma}^{(\mathcal{T},l)}=\frac{p}{\operatorname{tr}(\overline{\Sigma}^{(\mathcal{T},l)})}\overline{\Sigma}^{(\mathcal{T},l)}\enskip\text{and}\enskip\widehat{\Omega}^{(\mathcal{T},l)}=\frac{\operatorname{tr}(\overline{\Sigma}^{(\mathcal{T},l)})}{p}\overline{\Omega}^{(\mathcal{T},l)},$ (13)
@@
where $\overline{\Sigma}^{(\mathcal{T},l)}:=\overline{\Omega}^{(\mathcal{T},l)-1}$.

### v2

@@ 3.2.1 Single Edge Test
By leveraging information across $m$ sample groups using group Lasso (Eq. 2), we not only improve the efficiency of spatial partial correlation estimates but also enhance the power of tests for detecting significant associations. We first consider testing a single edge: for a pair $(i,j)$, the null hypothesis is
@@
Given the group sparsity structure, we construct a test statistic by summing the partial correlation estimates $\widehat{\rho}^{(\mathcal{S},l)}_{ij}$ across all sample groups. To account for varying sample sizes across sample groups, we weight the estimates by $\sqrt{n_{l}p}$ and define the test statistic as $\widehat{T}_{ij}:=\frac{1}{\sqrt{m}}\sum_{l=1}^{m}\sqrt{n_{l}p}\,\widehat{\rho}^{(\mathcal{S},l)}_{ij}$. In Proposition 4.2, we show that $\widehat{T}_{ij}$ asymptotically follows a normal distribution with mean $T_{ij}:=\frac{1}{\sqrt{m}}\sum_{l=1}^{m}\sqrt{n_{l}p}\,\rho^{(\mathcal{S},l)}_{ij}$. Assuming the sign of the associations remains consistent across sample groups, the mean equals zero if and only if the null hypothesis holds, so $\widehat{T}_{ij}$ should be significantly different from zero under the alternative hypothesis. Once the asymptotic variance is consistently estimated, constructing a confidence interval for $T_{ij}$ and the corresponding $p$-value for $H_{0,ij}$ is straightforward.
@@
More generally, additional sign information on $\rho^{(\mathcal{S},l)}_{ij}$ may be available. With this additional knowledge, we present a test statistic based on a linear combination of those $\widehat{\rho}^{(\mathcal{S},l)}_{ij}$ for $l\in[m]$, which is closely related to its $\ell_{1}$ norm. More specifically, with an edge-specific sign vector $\sigma_{ij}:=(\sigma^{(1)}_{ij},\dots,\sigma^{(m)}_{ij})^{\top}\in\{-1,1\}^{m}$, we replace $\widehat{T}_{ij}$ with the following sign-addressed test statistic $\widehat{T}_{ij,\sigma}:=\frac{1}{\sqrt{m}}\sum_{l=1}^{m}\sigma^{(l)}_{ij}\sqrt{n_{l}p}\widehat{\rho}^{(\mathcal{S},l)}_{ij}$. The normal approximation we establish in Section 4 also applies to the sign-addressed test statistic.
@@ 3.2.2 Simultaneous Test
Now we shift our focus to a more challenging problem of multiple edge testing. In this scenario, we aim to test whether there are no edges at all in a user-specified edge set $E\subseteq[p]\times[p]$, regardless of its size, which corresponds to the following null hypothesis:
@@
When $E$ consists of a single edge $(i,j)$, $H_{0,E}$ reduces to the single-edge null hypothesis $H_{0,ij}$ in Eq. 5. While multiple testing techniques, such as Bonferroni correction, can be applied to extend single-edge tests to multiple-edge tests, these methods are often overly conservative. Moreover, in neuroscience applications, the edge set $E$ typically represents connections between different brain areas, and its cardinality can grow on the order of $q^{2}$. As a result, even if we establish the asymptotic normality of a single $\widehat{T}_{ij}$, traditional multiple testing methods may not be valid in high-dimensional settings where $\lvert E\rvert$ grows rapidly.
@@
The key idea leverages on high-dimensional central limit theorems (e.g., Chernozhukov et al., 2013): Although the full vector $\widehat{T}_{E}$ may not be asymptotically jointly normal as $\lvert E\rvert$ increases, the supremum norm $\lVert\widehat{T}_{E}-T_{E}\rVert_{\infty}$ exhibits the same limiting behavior as $\lVert Z\rVert_{\infty}$, where $Z$ is a centered normal random vector with the same covariance as the asymptotic covariance of $\widehat{T}_{E}$. Specifically, this covariance is given by the matrix of asymptotic covariances between $\widehat{T}_{i_{1}j_{1}}$ and $\widehat{T}_{i_{2}j_{2}}$ for $(i_{1},j_{1}),(i_{2},j_{2})\in E$, expressed as (see Appendix E for the derivation):
[…]
We approximate the distribution of $\lVert Z\rVert_{\infty}$ using a parametric bootstrap based on the plug-in estimator $\widehat{S}_{EE}$ of the asymptotic covariance. Generating bootstrap samples $\{\widehat{Z}^{(1)},\dots,\widehat{Z}^{(B)}\}$ from $\operatorname{N}(0,\widehat{S}_{EE})$, we construct a $(1-\alpha)$-confidence region for $T_{E}$:
$\mathcal{C}_{E}(1-\alpha):=\{t\in\mathbb{R}^{\lvert E\rvert}:\lVert t-\widehat{T}_{E}\rVert_{\infty}\leq\widehat{q}_{1-\alpha}\},$ (8)
where $\widehat{q}_{1-\alpha}$ is the empirical $(1-\alpha)$-quantiles of $\{\lVert\widehat{Z}^{(1)}\rVert_{\infty},\dots,\lVert\widehat{Z}^{(B)}\rVert_{\infty}\}$. The null hypothesis $H_{0,E}$ is rejected if $\mathbf{0}\notin\mathcal{C}_{E}(1-\alpha)$. The coverage of this confidence region is studied in Theorem 4.3, with a power analysis provided in Theorem 4.4. We further discuss how borrowing information across sample groups enhances testing power in Remark 4.5.
@@ 3.3 Estimation of Temporal Covariance Matrix
The plug-in estimator $\widehat{S}_{EE}$ of the asymptotic covariance (Eq. 7) requires the Frobenius norm of the temporal covariance matrix, $\Sigma^{(\mathcal{T},l)}$ for each sample group. We propose to estimate $\Sigma^{(\mathcal{T},l)}$ based on a modified Cholesky decomposition of its inverse, the temporal precision matrix $\Omega^{(\mathcal{T},l)}$ (Bickel and Levina, 2008; Liu and Ren, 2020). The Cholesky decomposition characterizes the autoregressive dependence of the signal at each time point on its past values. Specifically, suppose for each sample group $l\in[m]$ that the modified Cholesky decomposition of $\Omega^{(\mathcal{T},l)}$ is $\Omega^{(\mathcal{T},l)}=L^{(\mathcal{T},l)}D^{(\mathcal{T},l)}L^{(\mathcal{T},l)\top},$ where $D^{(\mathcal{T},l)}$ is a diagonal matrix, and $L^{(\mathcal{T},l)}$ is a lower triangular matrix with diagonal entries equal to $1$. Let $\beta^{(\mathcal{T},l)}:=I-L^{(\mathcal{T},l)\top}$. Then the first (t1)(t-1) entries of $\beta^{(\mathcal{T},l)}$’s $t$-th row, $\beta^{(\mathcal{T},l)}_{\cdot,t}$, correspond to the regression coefficients of $X^{(k,l)}_{ti}$ on its past values, $X^{(k,l)}_{1i},X^{(k,l)}_{2i},\ldots,X^{(k,l)}_{(t-1)i}$ at each spatial location $i\in[q]$. That is, $X^{(k,l)}_{ti}=X^{(k,l)\top}_{\cdot,i}\beta^{(\mathcal{T},l)}_{\cdot,t}+\epsilon^{(\mathcal{T},k,l)}_{ti}$ for each time $t$. Due to the Kronecker product structure, the regression coefficients are invariant across spatial points $i\in[q]$, whereas the residuals exhibit spatial dependence characterized by $\Sigma^{(\mathcal{S},l)}$. To accommodate the matrix-variate data and the identifiability constraint $\operatorname{tr}(\Sigma^{(\mathcal{T},l)})=p$, we rewrite the Cholesky decomposition of $\Omega^{(\mathcal{T},l)}$ by
$\Omega^{(\mathcal{T},l)}=\frac{\operatorname{tr}(\Sigma^{(\mathcal{S},l)})}{q}(I-\beta^{(\mathcal{T},l)})^{\top}(\Phi^{(\mathcal{T},l)})^{-1}(I-\beta^{(\mathcal{T},l)}),$ (9)
where $\Phi^{(\mathcal{T},l)}:=\frac{\operatorname{tr}(\Sigma^{(\mathcal{S},l)})}{q}(D^{(\mathcal{T},l)})^{-1}$.
@@
In our working example, neural recordings are not guaranteed to be temporally aligned across sample groups due to response latencies to stimuli (Ventura, 2004). Hence we do not assume that $\Sigma^{(\mathcal{T},l)}$ is the same or even similar across sample groups $l\in[m]$. Instead, we impose a weaker bandable assumption on the modified Cholesky decomposition for each sample group. In physiological time-series signals, the dependence between time points naturally decays as the time lag increases, so a reasonable assumption is that $\beta^{(\mathcal{T},l)}_{st}$ approaches zero as $t-s\rightarrow\infty$. By applying bandable assumptions, we ignore weak dependencies between distant time points, achieving a bias-variance trade-off in the estimation of $\beta^{(\mathcal{T},l)}$.
@@
We estimate $\beta^{(\mathcal{T},l)}$ following the procedure in Liu and Ren (2020). Specifically, we treat the observed data $X^{(k,l)}$ as $q$ vector-variate “column samples”, $X_{\cdot,1}^{(k,l)},\dots,X_{\cdot,q}^{(k,l)}$ rather than as $p$ vector-variate “row samples” as in Section 3.1 for spatial covariance matrix estimation, and fit the following linear regression model at each time point $t$:
@@
$\widehat{\beta}^{(\mathcal{T},l)}_{\cdot,t}=\operatorname*{\arg\!\min}_{b\in\mathbb{R}^{p}}\frac{1}{2n_{l}q}\sum_{k=1}^{n_{l}}\lVert X^{(k,l)}_{t,\cdot}-X^{(k,l)\top}b\rVert_{2}^{2}$ (10)
@@
with respect to $b_{s}=0$ where $s<t-h_{l}$ or $s\geq t$, where the bandwidth hyperparameter $h_{l}$ can be either user-specified or data-driven. Unlike the spatial case, the estimation of the temporal covariance matrix is performed individually for each sample group. Having the regression coefficients estimated, we estimate $\Phi^{(\mathcal{T},l)}$ by
@@
$\widehat{\Phi}^{(\mathcal{T},l)}_{tt}:=\frac{1}{n_{l}q}\sum_{k=1}^{n_{l}}\left\lVert X^{(k,l)}_{t,\cdot}-X^{(k,l)\top}\beta^{(\mathcal{T},l)}_{\cdot,t}\right\rVert_{2}^{2}.$ (11)
@@
For technical issues, we truncate the eigenvalues of $I-\beta^{(\mathcal{T},l)}$ to ensure its invertibility, following the approach in Liu and Ren (2020). Given a cut-off hyperparameter $\eta>0$, for a square matrix $A$, let $P_{\eta}(A):=U~\max\{\min\{\Lambda,\eta\},\eta^{-1}\}~V^{\top}$, where $A$ has a singular value decomposition $A=U\Lambda V^{\top}$, and the $\min$ and $\max$ above are element-wise operations. Let $\overline{\Omega}^{(\mathcal{T},l)}$ be the precursor estimator of $\Omega^{(\mathcal{T},l)}$ given by
$\overline{\Omega}^{(\mathcal{T},l)}:=P_{\eta}(I-\widehat{\beta}^{(\mathcal{T},l)})^{\top}(\widehat{\Phi}^{(\mathcal{T},l)})^{-1}P_{\eta}(I-\widehat{\beta}^{(\mathcal{T},l)}).$ (12)
@@
Conforming to the identifiability constraint $\operatorname{tr}(\Sigma^{(\mathcal{T},l)})=p$, we propose our estimators of $\Sigma^{(\mathcal{T},l)}$ and $\Omega^{(\mathcal{T},l)}$ as
@@
$\widehat{\Sigma}^{(\mathcal{T},l)}:=\frac{p}{\operatorname{tr}(\overline{\Sigma}^{(\mathcal{T},l)})}\overline{\Sigma}^{(\mathcal{T},l)}\enskip\text{and}\enskip\widehat{\Omega}^{(\mathcal{T},l)}:=(\widehat{\Sigma}^{(\mathcal{T},l)})^{-1}=\frac{\operatorname{tr}(\overline{\Sigma}^{(\mathcal{T},l)})}{p}\overline{\Omega}^{(\mathcal{T},l)},$ (13)
@@
where $\overline{\Sigma}^{(\mathcal{T},l)}:=(\overline{\Omega}^{(\mathcal{T},l)})^{-1}$. The estimation error bounds of the Frobenius norm of $\Sigma^{(\mathcal{T},l)}$ are provided in Section C.2.